OpenStudy (anonymous):

Which of the following values is the 75th percentile? lower quartile median upper quartile

OpenStudy (campbell_st):

quartiles divide a data set into 4 equal parts Lower Quartile = 1/4 median = 1/2 Upper Quartile = ?

OpenStudy (anonymous):

i have no clue

OpenStudy (campbell_st):

well upper quartile = 3/4 so convert the fractions to percentages... and you'll find the answer

OpenStudy (anonymous):

its upper

OpenStudy (anonymous):

right? @campbell_st

OpenStudy (campbell_st):

correct